### 12 Dec 2023 v1

@@ 1 Introduction
Asking whether the Sun, or any star, has a central black hole is equivalently a question about dark matter. While the evidence for its existence is now overwhelming, the fundamental nature of dark matter has eluded explanation. It is clear from observations of big bang nucleosynthesis, the cosmic microwave background, large scale structure, cluster lensing, and galaxy rotation curves in spiral galaxies and velocity dispersions in elliptical galaxies, among a long list of other observations, that the majority of the matter in the universe simply does not interact electromagnetically [for a review see 35]. Notably, these observations are independent of one another and span cosmic time, probing the dark matter abundance from the first minutes of the universe to the present, and are all consistent in finding that 85% of the mass in the universe is invisible and thus is dark.
The composition of this invisible matter is unknown and candidates are too numerous to enumerate here, but can be broadly grouped based on their masses. Microscopic theories of dark matter generally consist of new particles beyond the Standard Model [see sec. 27.5 in 36]. In contrast macroscopic candidates may be composites of new particles from an entire dark sector, but many candidates only require Standard Model physics such as quark nuggets, strangelets, and most notably primordial black holes [2, 34, 25].
@@
This work is a primer on Hawking stars and is both a companion and extension of our recent work, [9]. We present a brief description of our implementation in the stellar evolution code Mesa in Sec. 2, and some basic theory of black hole accretion in Hawking stars in Sec. 3. In Sec. 3.3 we present stellar evolution models for a range of seed masses using the accretion scheme of [9]. In Sec. 3.4, we present a new accretion model developed for this work that accounts for photon trapping, along with detailed stellar evolution models. We summarize with some open questions in Sec. 4.
@@ 3 Black Hole Accretion in the Solar Core
Figure 1: (Top) luminosity and (bottom) accretion rate of a low mass black hole accreting from the solar core, assuming a constant radiative efficiency of $\epsilon=0.08$. For low black hole masses the accretion is limited by the convective luminosity (solid black) while at higher accretion rates the radiation pressure from the accretion luminosity limits the accretion to an Eddington rate (dashed black); they intersect at the Bondi-Eddington transition (BET), $M_{\rm BET}=2\times 10^{-10}$ $\textrm{M}_{\odot}$. The actual evolutionary path is shaded. At no point does this model accrete at the sound speed limited Bondi rate (red). This may be a useful toy model for a rapidly rotating or magnetized star, and is implemented in the stellar evolution calculations shown in Fig. 2.
@@ 3.2 Convection-Limited Bondi Accretion
By equating Eqs. 1 and 5 one can find a differential equation of the same form as Eq. 2. While this equation still diverges at a finite time, it does so at
@@ 3.3 Eddington Accretion
where $\kappa$ is the opacity of the solar core ($\kappa=1.5~{}\textrm{cm}^{2}~{}\textrm{g}^{-1}$). The accretion rate transitions from a Bondi rate to an Eddington rate approximately when Eqs. 5 and 7 are equal. The black hole mass at the Bondi-Eddington transition (BET) for the solar core is approximately $M_{\mathrm{BET}}\approx 2\times 10^{-10}$ $\textrm{M}_{\odot}$ with an accretion luminosity of $10^{-5}$ $\textrm{L}_{\odot}$. More generally, the BET depends on both the metallicity and mass of the star as these set $\kappa$, $c_{s}$, and $\rho$. We emphasize that this is only a rough first approximation, and the core gas pressure can also play an important role in regulating the accretion rate and should be studied using more detailed accretion schemes like those of [22, 23].
When combining Eq. 7 and 1, we obtain a differential equation where $\dot{M}_{\mathrm{BH}}\propto M_{\mathrm{BH}}$, so the black hole grows exponentially as $M_{\mathrm{BH}}(t)=M_{\mathrm{PBH}}e^{t/\tau_{E}}$ with a growth timescale of
@@
In Fig. 2 we show three examples for the evolution of an initially Sun-like Hawking star with a seed PBH in the core calculated with Mesa. For seed masses less than about $10^{-10}$ $\textrm{M}_{\odot}$, the appearance of the Sun is largely unchanged at present age of the Sun (4.6 Gyr) and the effects of the black hole are confined to the inner core. Eventually, the black hole grows large enough that the accretion luminosity dominates over fusion, causing the star to expand and thus quenching fusion. We consider this a useful marker for delimiting the ‘main sequence’ from the ‘post-main sequence.’
@@
The transition of a Hawking star from its post-main sequence phase to some other high energy transient is not resolved here, as the Mesa models fail when the Bondi radius reaches the surface of the star. At this time the black hole mass is roughly $10^{-6}$ $\textrm{M}_{\odot}$, or roughly an earth mass. Significantly more work is required to bridge the late stages of evolution in simulations.
@@ 3.4 Photon Trapping and Super-Eddington Accretion
Figure 4: (Top) Kippenhahn diagrams of late stage evolution of a central black hole for a Sun-like Hawking star including photon trapping. Colors and shading is the same as 1. Early times, prior to the transition to an Eddington-like accretion, are largely indistinguishable from the models with constant $\epsilon$ in Fig. 1. The break when about a thousand years remain is associated with the onset of trapping, while the final break with a few hours remaining is associated with a rapid increase in nuclear burning. The time until destruction is taken to be the time until the end of the simulation. (Bottom) The accretion rapidly accelerates as the accretion re-enters the Bondi regime, and the mass grows rapidly to about $10^{-2}$ $\textrm{M}_{\odot}$ before the simulation ends for numerical reasons.
In Fig. 4 we show two stellar evolution models that depict the rapid late time growth of the central black hole with photon trapping for initial $10^{-11}$ $\textrm{M}_{\odot}$  and $10^{-12}$ $\textrm{M}_{\odot}$ seeds. For numerical reasons both the radiative efficiency and accretion luminosity have a floor of $1.1\times 10^{-20}$. At early times when $\epsilon\approx 0.08$, the model with a adaptive radiative efficiency is largely indistinguishable from the first model, and for sufficiently low seed masses ($M_{\mathrm{BH}}\lesssim 10^{-12}$ $\textrm{M}_{\odot}$) the star survives for its entire main sequence and accumulates a large He core. However, as a consequence of this new accretion scheme Sun-like Hawking stars do not have an extended post-main sequence evolution.
@@ 4 Summary
The two models presented in this work may be best thought of as extremes, showing how a high radiative efficiency or a low radiative efficiency can drastically influence the evolution toward a long-lived sub-subgiant phase or a rapid direct collapse. It is not immediately clear how suitable either of these models are for describing the real physics of microscopic black hole accretion in a main sequence star in the modern universe. One might reasonably guess that a star could evolve following one of these two qualitative evolutionary tracks depending on the stellar mass, metallicity, rotation, and magnetic field. In this work, we have only explored one fiducial star and a few seed masses using very simplified accretion schemes. Future theory work will be needed to understand accretion in Hawking stars and implement these accretion schemes in stellar evolution codes. Fortunately, a large body of theory on microscopic black hole accretion already exists and could be implemented in Mesa in the near future [e.g. 5, 6, 33, 22, 23, 29, 20].
@@ Acknowledgments
The authors thank Taeho Ryu, Deepika Bollimpalli, Warrick Ball, Florian Kühnel, Selma de Mink, and Jørgen Christensen-Dalsgaard for their contributions to the work this article is based on. Financial support for this publication comes from Cottrell Scholar Award #CS-CSA-2023-139 sponsored by Research Corporation for Science Advancement. This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470) to MC.

### 19 Jan 2024 v2

@@ 1 Introduction
Asking whether the Sun, or any star, has a central black hole is equivalently a question about dark matter. While the evidence for its existence is now overwhelming, the fundamental nature of dark matter has eluded explanation. It is clear from observations of big bang nucleosynthesis, the cosmic microwave background, large scale structure, cluster lensing, and galaxy rotation curves in spiral galaxies and velocity dispersions in elliptical galaxies, among a long list of other observations, that the majority of the matter in the universe simply does not interact electromagnetically [for a review see 37]. Notably, these observations are independent of one another and span cosmic time, probing the dark matter abundance from the first minutes of the universe to the present, and are all consistent in finding that 85% of the mass in the universe is invisible and thus is dark.
The composition of this invisible matter is unknown and candidates are too numerous to enumerate here, but can be broadly grouped based on their masses. Microscopic theories of dark matter generally consist of new particles beyond the Standard Model [see sec. 27.5 in 38]. In contrast macroscopic candidates may be composites of new particles from an entire dark sector, but many candidates only require Standard Model physics such as quark nuggets, strangelets, and most notably primordial black holes [2, 36, 27].
@@
This work is a primer on Hawking stars and is both a companion and extension of our recent work, [9]. We begin with an estimate of PBH capture rates in Sec. 2. We present a brief description of our implementation in the stellar evolution code Mesa in Sec. 3, and some basic theory of black hole accretion in Hawking stars in Sec. 4. In Sec. 4.3 we present stellar evolution models for a range of seed masses using the accretion scheme of [9]. In Sec. 4.4, we present a new accretion model developed for this work that accounts for photon trapping, along with detailed stellar evolution models. We summarize with some open questions in Sec. 5.
2 Primordial Black Hole Capture Rate
In this section we briefly estimate the capture rates of PBHs by stars in the Milky Way using the methods of [21] and [22]. The mean number of asteroid-mass PBHs captured by a star is $\bar{N}=f\eta\,\nu(M)$ where $f$ is the PBH fraction of dark matter, $\eta$ is the ‘merit factor’ of the host galaxy (see below), and $\nu(M)$ is a factor of order unity that depends on the mass of the star. The factor $\nu(M)$ is determined numerically from simulations of contraction of star forming clouds by [21], and takes into account the amount of time required for the PBH to sink into the star. They show that $\nu(M)$ is largely independent of PBH mass between $10^{-15}$ $\textrm{M}_{\odot}$ and $10^{-13}$ $\textrm{M}_{\odot}$, but has some weak dependence above this mass. It was only determined for stars up to 0.8 $\textrm{M}_{\odot}$, so future work is required for more massive stars.
The merit factor depends on the dark matter density $\rho_{\rm{DM}}$ and velocity dispersion $\sigma$ of the halo, and dominates the capture probability. If PBHs constitute all the dark matter (${f=1}$) then the merit factor directly gives an order-of-magnitude probability that a star-forming cloud will capture a PBH. The merit factor is
$\eta=\left(\frac{\rho_{\rm{DM}}}{100\,\rm{GeV/cm}^{3}}\right)\left(\frac{7\,\mathrm{km/s}}{\sqrt{2}\sigma}\right)^{3}$ (1)
and assumes that only PBHs in the low-velocity tail are captured. For the Milky Way disk, $\eta\approx 10^{-7}$, so it is highly unlikely that the Sun is a Hawking star even if PBHs comprise all the dark matter. However, it does suggest there could be of order $10^{4}$ Hawking stars in the Galactic disk. In the Gaia Data Release 3 one might expect dozens to hundreds of Hawking stars [25]. In the sections that follow, we will show how these can be searched for.
Perhaps more importantly, [21] shows that most stars in ultra-faint dwarfs like Tucana III (${\eta=0.51}$) and Triangulum II (${\eta=0.95}$) may have captured PBHs upon formation. This has incredible potential for constraining PBHs in the asteroid mass window as dark matter, but that is only if the evolution and survival times and of Hawking stars are well understood.
We emphasize that more work is needed on PBH captures over a larger range of stellar masses and PBH masses in order to put strong constraints on PBH dark matter using Hawking stars, but these rough order of magnitude scales suffice for motivating this work.
@@ 4 Black Hole Accretion in the Solar Core
Figure 1: (Top) luminosity and (bottom) accretion rate of a low mass black hole accreting from the solar core, assuming a constant radiative efficiency of $\epsilon=0.08$. For low black hole masses the accretion is limited by the convective luminosity (solid black) while at higher accretion rates the radiation pressure from the accretion luminosity limits the accretion to an Eddington rate (dashed black); they intersect at the Bondi-Eddington transition (BET), $M_{\rm BET}=2\times 10^{-10}$ $\textrm{M}_{\odot}$. The actual evolutionary path is shaded. At no point does this model accrete at the sound speed limited Bondi rate (blue). This may be a useful toy model for a rapidly rotating or magnetized star, and is implemented in the stellar evolution calculations shown in Fig. 2.
@@ 4.2 Convection-Limited Bondi Accretion
By equating Eqs. 2 and 6 one can find a differential equation of the same form as Eq. 3. While this equation still diverges at a finite time, it does so at
@@ 4.3 Eddington Accretion
where $\kappa$ is the opacity of the solar core ($\kappa=1.5~{}\textrm{cm}^{2}~{}\textrm{g}^{-1}$). The accretion rate transitions from a Bondi rate to an Eddington rate approximately when Eqs. 6 and 8 are equal. The black hole mass at the Bondi-Eddington transition (BET) for the solar core is approximately $M_{\mathrm{BET}}\approx 2\times 10^{-10}$ $\textrm{M}_{\odot}$ with an accretion luminosity of $10^{-5}$ $\textrm{L}_{\odot}$. More generally, the BET depends on both the metallicity and mass of the star as these set $\kappa$, $c_{s}$, and $\rho$. We emphasize that this is only a rough first approximation, and the core gas pressure can also play an important role in regulating the accretion rate and should be studied using more detailed accretion schemes like those of [23, 24].
When combining Eq. 8 and 2, we obtain a differential equation where $\dot{M}_{\mathrm{BH}}\propto M_{\mathrm{BH}}$, so the black hole grows exponentially as $M_{\mathrm{BH}}(t)=M_{\mathrm{PBH}}e^{t/\tau_{E}}$ with a growth timescale of
@@
In Fig. 2 we show three examples of the evolution of an initially Sun-like Hawking star with a seed PBH in the core calculated with Mesa. For seed masses less than about $10^{-10}$ $\textrm{M}_{\odot}$, the appearance of the Sun is largely unchanged at present age of the Sun (4.6 Gyr) and the effects of the black hole are confined to the inner core. Eventually, the black hole grows large enough that the accretion luminosity dominates over fusion, causing the star to expand and thus quenching fusion. At this time the black hole mass is about $10^{-5}$ $\textrm{M}_{\odot}$, or roughly a Uranus mass. We consider this a useful marker for delimiting the ‘main sequence’ from the ‘post-main sequence.’
@@
The transition of a Hawking star from its post-main sequence phase to some other high energy transient is not resolved here, as the Mesa models fail when the Bondi radius reaches the surface of the star ($M_{\textrm{BH}}\approx 10^{-1}$ $\textrm{M}_{\odot}$). Significantly more work is required to bridge the late stages of evolution in simulations.
@@ 4.4 Photon Trapping and Super-Eddington Accretion
Figure 4: (Top) Kippenhahn diagrams of late stage evolution of a central black hole for a Sun-like Hawking star including photon trapping. Colors and shading is the same as Fig. 1. Early times, prior to the transition to an Eddington-like accretion, are largely indistinguishable from the models with constant $\epsilon$ in Fig. 1. The break when about a thousand years remain is associated with the onset of trapping, while the final break with a few hours remaining is associated with a rapid increase in nuclear burning. The time until destruction is taken to be the time until the end of the simulation. (Bottom) The accretion rapidly accelerates as the accretion re-enters the Bondi regime, and the mass grows rapidly to about $10^{-2}$ $\textrm{M}_{\odot}$ before the simulation ends for numerical reasons.
In Fig. 4 we show two stellar evolution models that depict the rapid late time growth of the central black hole with photon trapping for initial $10^{-11}$ $\textrm{M}_{\odot}$  and $10^{-12}$ $\textrm{M}_{\odot}$ seeds. For numerical reasons both the radiative efficiency and accretion luminosity have a floor of $1.1\times 10^{-20}$. At early times when $\epsilon\approx 0.08$, the model with an adaptive radiative efficiency is largely indistinguishable from the first model, and for sufficiently low seed masses ($M_{\mathrm{BH}}\lesssim 10^{-12}$ $\textrm{M}_{\odot}$) the star survives for its entire main sequence and accumulates a large He core. However, as a consequence of this new accretion scheme Sun-like Hawking stars do not have an extended post-main sequence evolution.
@@ 5 Summary
The two models presented in this work may be best thought of as extremes, showing how a high radiative efficiency or a low radiative efficiency can drastically influence the evolution toward a long-lived sub-subgiant phase or a rapid direct collapse. It is not immediately clear how suitable either of these models are for describing the real physics of microscopic black hole accretion in a main sequence star in the modern universe. One might reasonably guess that a star could evolve following one of these two qualitative evolutionary tracks depending on the stellar mass, metallicity, rotation, and magnetic field. In this work, we have only explored one fiducial star and a few seed masses using very simplified accretion schemes. Future theory work will be needed to understand accretion in Hawking stars and implement these accretion schemes in stellar evolution codes. Fortunately, a large body of theory on microscopic black hole accretion already exists and could be implemented in Mesa in the near future [e.g. 5, 6, 35, 23, 24, 31, 20].
@@ Acknowledgments
The authors thank Taeho Ryu, Deepika Bollimpalli, Warrick Ball, Florian Kühnel, Selma de Mink, and Jørgen Christensen-Dalsgaard for their contributions to the work this article is based on, and Nicolas Esser for discussion. Financial support for this publication comes from Cottrell Scholar Award #CS-CSA-2023-139 sponsored by Research Corporation for Science Advancement. This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470) to MC.